OpenStudy (anonymous):

f(x)=2x^2-x-1. If f(x)=-1, what is x?

jimthompson5910 (jim_thompson5910):

f(x)=2x^2-x-1 -1=2x^2-x-1 ... replace f(x) with -1 (since f(x) = -1) -1+1=2x^2-x-1+1 0 = 2x^2 - x 2x^2 - x = 0 I'll let you finish

OpenStudy (anonymous):

That's what I have right now. I don't know what to do next. I factored out an x so I have\[x(2x-1)=0\]

jimthompson5910 (jim_thompson5910):

remember that if A*B = 0, then either A = 0 or B = 0 (or both could equal zero)

jimthompson5910 (jim_thompson5910):

so this means that if x*(2x-1) = 0 then x = 0 or 2x - 1 = 0

OpenStudy (anonymous):

I keep forgetting basic stuff... Thanks.

jimthompson5910 (jim_thompson5910):

yw